What is the value or values in the equation p2+13p-30=0
Answer by ewatrrr(24785)


Hi,
solving
p^2+13p-30=0
(p+15)(p-2)= 0 Note:SUM of the inner product(15p) and the outer product(-2p) = 13p
(p+15)= 0
p = -15
(p-2)= 0
p = 2